# MEIT: Multi-Modal Electrocardiogram Instruction Tuning on Large Language Models for Report Generation

### v3

@@ 5.3 Analysis
6 Related Works
Instruction Tuning. Instruction tuning [36, 37] boosts zero-shot learning in LLMs for new tasks using instructions. Notable models like InstructGPT [35], FLAN-PaLM [38],
and Alpaca [39] fine-tune with instruction data through various methods, including human feedback. Similarly, multimodal models such as LLaVA [33], MiniGPT-4 [40], and AnyMAL [41] benefit from multimodal instructions for enhanced learning. However, instruction tuning for medical signals like ECG remains underexplored. Our research develops a targeted instruction-tuning framework for ECG report generation, addressing this gap.
Medical Report Generation. Medical report generation [42]involves three strategies: (1) Template Selection and Generation, highlighted by HRGR [43] and CMAS [44]; (2) Data Integration and Coherence, as seen in PPKED [45] and CA [46]; (3) Cross-Modal Alignment, with efforts like [47, 48]. While recent research has advanced these areas, applying them to ECG data is challenging due to the distinct nature of ECG signals compared to medical images.
LLM for ECG Understanding. Only a few research efforts have focused on utilizing LLMs for ECG-related tasks [49, 50, 51]. In the studies by [49, 51], the authors manually convert ECG signals into text features before feeding them into LLMs, rather than using the original ECG data. This approach overlooks the modality patterns inherent in the signals. Moreover, these studies only employ LLMs for classifying diseases from ECG data and do not explore the generation of medical reports. In contrast, [50] attempts ECG report generation by using handcrafted ECG features as inputs instead of the original signals. However, they have not publicly made any code or models available, making direct comparisons challenging.
7 Conclusion and Future Work
In this work, we propose and demonstrate the efficacy of the Multimodal ECG Instruction Tuning (MEIT) framework. Particularly, we first introduce a novel method for creating instruction-following data for ECG reports, enabling the training of an ECG report generator, a multimodal LLM designed to generate ECG reports following human instructions. Furthermore, we propose an effective and efficient method for aligning ECG and report representations across multiple open-source LLMs, achieving notable performance on both the MIMIC-IV-ECG and PTB-XL datasets for diverse tasks. Importantly, this paper introduces the comprehensive benchmark for ECG instruction-following in ECG report generation. While this research primarily concentrates on ECG signals, it marks a preliminary step in biomedical signal instruction tuning. In future work, we plan to extend our framework to encompass additional medical signal domains, such as EEG signals, with the aspiration that our work spurs further advancements in developing more proficient medical-signal LLMs.
Appendix A Appendix.
A.1 Limitations and Future Work
In this work, we propose an effective and efficient method for generating ECG reports. Since this is a generative task, the content produced may occasionally lead to hallucinations. This is particularly pertinent in medical contexts such as ECG report generation, where accuracy is of utmost importance. To mitigate this issue, future iterations of our work could incorporate retrieval-augmented methods [14], information optimized strategies [52, 53, 54]. For instance, by integrating a retrieval mechanism that accesses a database of verified medical or biological knowledge [55] or previous ECG reports, the model could be guided towards generating more accurate and reliable outputs.
@@ A.2 Hyper-parameters of ECG Instruction Tuning
In this study, we implement the Low-Rank Adaptation (LoRA) [16] technique for efficient fine-tuning, specifically applied to ECG instruction tuning. As detailed in Table A.2 provided, we utilize mixed precision at bf16 for enhanced computational efficiency. Our models undergo instruction tuning over 5 epochs, with LoRA parameters set at an alpha of 64 and a rank of 128, accompanied by a dropout rate of 0.1. The total batch size is 64, with a gradient accumulation factor of 2. The maximum sequence length is constrained to 256 tokens. Additionally, we adopt a learning rate with 2e-5 for GPT-NeoX and 1e-4 for the other models, optimized using the AdamW algorithm. The learning rate follows a linear schedule with a warm-up ratio of 0.03. We set the weight decay to 0.0.
Moreover, as shown in Table A.2, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding. GPT2-Medium and GPT2-Large feature ECG dimensions 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128. GPT-NeoX employs a dimension of 96, and GPT-J notably uses the largest dimension of 256. These dimensions, reflecting each model’s head dimension design, illustrate diverse strategies in ECG data processing across different models.
@@ A.3 More Details of ECG Encoder
Projection Layer For the design of the projection layer within the ECG encoder, we adopt a non-linear approach similar to CLIP [56] and Med-UniC [57]. Specifically, in our experiments, we employ two consecutive linear layers, each followed by BatchNorm1d [^4] https://pytorch.org/docs/stable/generated/torch.nn.BatchNorm1d.html. Besides, ReLU serves as the activation function between the two linear layers. The default settings for input and hidden layers dimensions are set to 2048 in our experiment.
@@ A.4 Further Analysis of Generated Prompts
In the ECG instruction data curation, we manually created 32 prompt examples, as illustrated in Section 3.1. To increase the diversity of our samples, we employed GPT-4 to rephrase these manually designed prompts, generating a larger pool of prompt examples. These generated examples were randomly sampled and paired with ECG-text pairs to compile the ECG instruction dataset. In this section, We compare the experiment’s effects using 128, 256, and 512-generated samples, respectively. Table 8 shows the corresponding results with different dimensions. When the number is 256, it can achieve better results in most experimental settings. Hence, we take 256 generated samples as our default setting during the instruction tuning and inference.
@@
A.5 Visualization of Generated ECG Report Samples
As illustrated in Figures 9,  10, and  11 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct. The samples are presented in blue font to highlight the key information that aligns with the ground truth. The visualization demonstrates that all three selected models can capture the essential patterns of ECG signals and generate accurate reports. This underscores the efficacy of our proposed MEIT framework, which is adaptable to most open-source LLMs. It effectively learns the correct clinical semantics of ECG signals, thereby enabling the generation of corresponding reports.

### v4

@@ A.2 Hyper-parameters of ECG Instruction Tuning
In this study, we implement the Low-Rank Adaptation (LoRA) (Hu et al., 2021) technique for efficient fine-tuning, specifically applied to ECG instruction tuning. As detailed in Table 8 provided, we utilize mixed precision at bf16 for enhanced computational efficiency. Our models undergo instruction tuning over 5 epochs, with LoRA parameters set at an alpha of 64 and a rank of 128, accompanied by a dropout rate of 0.1. The total batch size is 64, with a gradient accumulation factor of 2. The maximum sequence length is constrained to 256 tokens. Additionally, we adopt a learning rate with 2e-5 for GPT-NeoX and 1e-4 for the other models, optimized using the AdamW algorithm. The learning rate follows a linear schedule with a warm-up ratio of 0.03. We set the weight decay to 0.0.
Moreover, as shown in Table 8, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding. GPT2-Medium and GPT2-Large feature ECG dimensions 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128. GPT-NeoX employs a dimension of 96, and GPT-J notably uses the largest dimension of 256. These dimensions, reflecting each model’s head dimension design, illustrate diverse strategies in ECG data processing across different models.
@@ A.3 More Details of ECG Encoder
Projection Layer For the design of the projection layer within the ECG encoder, we adopt a non-linear approach similar to CLIP (Radford et al., 2021) and Med-UniC (Wan et al., 2024a). Specifically, in our experiments, we employ two consecutive linear layers, each followed by BatchNorm1d [^4] https://pytorch.org/docs/torch.nn.BatchNorm1d.html. Besides, ReLU serves as the activation function between the two linear layers. The default settings for input and hidden layers dimensions are set to 2048 in our experiment.
@@
Table 9: Comparisons of results with and without supervised manner. We take LLaMA-2-Instruct as the LLM backbone here. heavy teal color indicates the highest results.
Methods Size MIMC-IV-ECG PTB-XL
BLEU-4 METEOR ROUGE-L CIDEr-D MTA MTA LC DA
Vision Mamba 86M 0.548 0.737 0.715 5.58 3.78 3.88 3.61 3.50
Vision Transformer 98M 0.592 0.815 0.772 5.67 4.33 4.15 4.12 3.78
Vision Transformer (SSL) 98M 0.581 0.822 0.766 5.75 4.42 4.28 3.85 3.85
1-D Temporal Conv (Ours) 20.4M 0.610 0.799 0.773 5.78 4.52 4.38 4.01 3.98
Ablation Study of ECG Encoder we conducted additional experiments comparing our default 1-D Temporal Convolution ECG encoder with alternative architectures, including: 1. S4-based Model: Vim-B (Vision Mamba, 98M parameters) (Zhu et al., 2024a). 2. Transformer-based Model: ViT-B/16 (Vision Transformer, 86M parameters) (Dosovitskiy et al., 2020), adapted for 1-D token patching to align with the temporal nature of ECG signals. 3. SSL-Transformer Model: ViT-B/75 initialized with self-supervised learning (SSL) weights specific to ECG signals (Na et al., 2024). We evaluated these models on two tasks: Quality of Generated Reports using the MIMIC-IV-ECG dataset, and Evaluation of Alignment with Human Expert Annotations using the PTB-XL dataset. For fair comparison, we used Meta-Llama-3-8B-Instruct as the LLM backbone due to its consistent strong performance.
The results, summarized in the table below, show that our 1-D Temporal Convolution ECG encoder, despite having significantly fewer parameters, performs comparably or better across most metrics compared to ViT and ViT-SSL, and comprehensively outperforms the S4-based Vim. Notably, the ViT-SSL encoder demonstrates the benefit of self-supervised pretraining for initial ECG representation learning. However, our default ECG encoder effectively captures the 12-channel ECG temporal patterns while remaining lightweight, making it well-suited for our efficient instruction tuning framework. These findings validate the effectiveness of our 1-D Temporal Conv encoder and also provide valuable insights for future work, including designing more complex ViT-based architectures optimized for ECG time-series data.
@@ A.4 Further Analysis of Generated Prompts
Prompts Number Analysis In the ECG instruction data curation, we manually created 32 prompt examples, as illustrated in Section 3.3. To increase the diversity of our samples, we employed GPT-4 to rephrase these manually designed prompts, generating a larger pool of prompt examples. These generated examples were randomly sampled and paired with ECG-text pairs to compile the ECG instruction dataset. In this section, We compare the experiment’s effects using 128, 256, and 512-generated samples, respectively. Table 11 shows the corresponding results with different dimensions. When the number is 256, it can achieve better results in most experimental settings. Hence, we take 256 generated samples as our default setting during the instruction tuning and inference.
@@
Ablation Study on GPT-4 Prompt Rephrasing We also conducted an ablation study to compare the performance with and without GPT-4 rephrasing prompts, using a fixed prompt for the latter. The results in the following Table 12 indicate that using diverse prompts rephrased by GPT-4 leads to better performance, highlighting the superiority of instruction tuning in enhancing LLMs’ generalization to new tasks and data over direct fine-tuning.
Table 12: Performance comparison of with and without GPT-4 prompt rephrasing. We take Mistral-Instruct as the LLM backbone here. heavy teal color indicates the highest results.
Prompt Nums BLEU-4 METEOR ROUGE-L CIDEr-D
w.o. Rephrasing 0.564 0.745 0.738 5.50
w. Rephrasing (Ours) 0.576 0.768 0.751 5.62
A.5 Comparison with Encoder-Decoder Models
In this section, we conducted additional comparative experiments using two open-source traditional encoder-decoder architectures: BART-Large (406M parameters) (Lewis, 2019) and T5-Large (780M parameters) (Raffel et al., 2020), as shown in Table 13. In adapting our framework for ECG instruction tuning, we employ the language encoder to process the input instruction, an ECG encoder to handle the input ECG signals, and the language decoder to generate the ECG report based on the output from both language end ECG encoder.
Our findings indicate that the performance of encoder-decoder models is comparable to the small pre-trained language models (GPT2-Medium and GPT-Large) presented in Table 1 and Table 2 of our paper. Moreover, LLM-based backbones (such as LLaMA1-2) consistently achieve a significant margin of improvement over the encoder-decoder architectures across all metrics.
Table 13: Comparison with encoder-decoder-based models on MIMIC-IV-ECG. For model size, ’M’ denotes the million level, and ’B’ denotes the billion level. The light teal color indicates the second highest results, and heavy teal color indicates the highest results.
BART-Large 406M 0.525 0.498 0.466 0.388 0.455 0.472 0.5124 0.451 3.15
T5-Large 780M 0.595 0.542 0.465 0.422 0.498 0.456 0.522 0.438 4.08
LLaMA-1 7B 0.685 0.648 0.615 0.543 0.761 0.724 0.742 0.642 5.26
LLaMA-2-Instruct 7B 0.706 0.662 0.622 0.581 0.775 0.745 0.768 0.664 5.55
A.6 Analysis of Combining MEIT with a Supervised Manner
In this section, we conduct a new experiment where we trained a CNN (ECG encoder) in a supervised manner on the PTB-XL training set, utilizing all available annotations (approximately 70 patterns), as shown in Table 14. We then transferred the CNN for ECG instruction fine-tuning on both the MIMIC-IV-ECG and PTB-XL datasets. Our findings indicate that performance increased on the PTB-XL dataset in most metrics, likely due to the model’s prior learning of specific annotated patterns. However, performance fluctuated on the MIMIC-IV-ECG dataset, which contains more data and exhibits greater diversity. This suggests that the supervised approach may enhance performance on in-domain data, but it limits generalizability to data from unseen domains.
Table 14: Comparisons of results with and without supervised manner. We take LLaMA-2-Instruct as the LLM backbone here. heavy teal color indicates the highest results.
Methods PTB-XL
BLEU-4 METEOR ROUGE-L CIDEr-D
MEIT 0.439 0.675 0.594 4.05
MEIT + Supervised manner 0.445 0.664 0.612 4.12
MIMIC-IV-ECG
BLEU-4 METEOR ROUGE-L CIDEr-D
MEIT 0.581 0.775 0.745 5.55
MEIT + Supervised manner 0.578 0.778 0.739 5.47
A.7 Computational Cost Analysis of MEIT
The time cost experiment, detailed in the Table 15, was conducted on the MIMIC-IV-ECG dataset. We found that larger models have longer training and inference times. Thus, we are considering techniques like quantization and other compression methods to improve model efficiency in future work.
Table 15: Computational time Analysis of MEIT with various parameters and backbones.
Model Size Training time Testing time
4 A100 and 3 Epochs 1 A100 and 128 Generated Samples
GPT-2 Large 774M 3.25h 3.125min
LLaMA-2-Instruct 7B 13.5h 9 min
LLaMA-2-Instruct (+) 13B 27h 14.125 min
A.8 Examples of ECG Instruction Tuning Data and the Corresponding Generated ECG Reports
As illustrated in Figures 8,  9, and  10 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct. The samples are presented in blue font to highlight the key information that aligns with the ground truth. The visualization demonstrates that all three selected models can capture the essential patterns of ECG signals and generate accurate reports. This underscores the efficacy of our proposed MEIT framework, which is adaptable to most open-source LLMs. It effectively learns the correct clinical semantics of ECG signals, thereby enabling the generation of corresponding reports.
@@ A.8 Examples of ECG Instruction Tuning Data and the Corresponding Generated ECG Reports
A.9 Future Works
In this work, we propose an effective and efficient method for generating ECG reports. Since this is a generative task, the content produced may occasionally lead to hallucinations Zhu et al. (2025). This is particularly pertinent in medical contexts such as ECG report generation, where accuracy is of utmost importance. To mitigate this issue, future iterations of our work could incorporate retrieval-augmented methods Wan et al. (2022); Li et al. (2024), information optimized strategies Xiong et al. (2022); Liang et al. (2020a, b), multi-modal fusion and optimization strategies Wan et al. (2024c); Xiong et al. (2024a); Liu et al. (2024c); Gong et al. (2024); Wan et al. (2023a); Chen et al. (2025); Luo et al. (2024); Shen et al. (2025b); Wan et al. (2025b); Zhang et al. (2025b); Liu et al. (2024b); Huang et al. (2024); Zhu et al. (2024b); Liu et al. (2025a), Reasoning enhanced methods Wan et al. (2025a); Liu et al. (2025b); Shen et al. (2025a), efficient ML Wang et al. (2024b); Tao et al. (2024); Wan et al. (2024b); Liu (2024); Xin et al. (2024); Shen et al. (2024); Wang et al. (2025); Xiong et al. (2024b, 2025). For instance, by integrating a retrieval mechanism that accesses a database of verified medical or biological knowledge Zheng et al. (2024) or previous ECG reports, the model could be guided towards generating more accurate and reliable outputs.